The perimeter of a rectangle is 144 inches. The length exceeds the width by 46 inches. Find the length and the width.
The length in inches ??
The width in inches ??
Help me please. I am so failing this clas s Ijust cant grasp this stuff. Help!!

Answer by scott8148(6628)

let x=width

"The length exceeds the width by 46 inches" __ so x+46=length

perimeter is twice the length plus twice the width __ 144=2(x+46)+2(x)

144=2x+92+2x __ 144=4x+92 __ 52=4x __ 13=x

width is 13 and length is 59